OpenStudy (anonymous):

an egg is dropped from the top of a fictitious tower which stands 1,111 feet tall. determine when the egg will hit the ground and splatter. - what formula do I use for this, and how do I use it.

OpenStudy (anonymous):

Use second law of motion: \[\huge h= ut + \frac{1}{2} gt^2\] Here s = 1111 ft, g= 9.8 m/s^2, u = 0 m/s, t =?

OpenStudy (anonymous):

sorry h= 1111ft and not s=1111